The resistance R of "\theta" is equal to Ro(1+"\alpha" "\theta" + "\beta" "\theta" 2). Where "\alpha" =3.8*103 and "\beta" = -5.6*10-7. What temparature will the platinium thermometer indicate when the temparature on the gas scale is 200K


1
Expert's answer
2021-12-15T11:19:58-0500

R=R0(1+"\alpha" "\theta" + "\beta" "\theta" 2),

"R_1=1.74R_0."